Consider four uniformly charged plates oriented normal to the x-axis as shown. The drawing scale exaggerates the size of the

lengths along x compared to sizes along y

and z. Let L=1.0m, d=1.0mm, Q1=3.0C,

Q2=-3.0C, Q3=8.0C, and Q4=-6.0C. The positions of the plates are at x1, x2, x3 and x4.

  1. a) Find the 2D charge density of each plate
  2. b) Findthenetelectricfieldinregion
  3. x2 <x<x3. 
1
Expert's answer
2021-03-10T17:14:50-0500

Plate area, A=l×d=1103=103m2A=l \times d = 1*10^{-3}= 10^{-3} m^2


Charge density of first plate, σ1=q1A=3103=3×103Cm2\sigma_1=\frac{q_1}{A}=\frac{3}{10^{-3}}=3\times 10^{-3} Cm^{-2}

Charge density on second plate, σ2=q2A=3103=3×103Cm2\sigma_2=\frac{q_2}{A}=\frac{-3}{10^{-3}}=-3\times 10^{-3} Cm^{-2}


Charge density on third plate, σ3=q3A=8103=8×103Cm2\sigma_3=\frac{q_3}{A}=\frac{8}{10^{-3}}=8\times 10^{-3} Cm^{-2}


Charge density on fourth plate, σ4=q4A=6103=6×103Cm2\sigma_4=\frac{q_4}{A}=\frac{-6}{10^{-3}}=-6\times 10^{-3} Cm{-2}


Let separation between the plates is 1m.

Electric field is given by, E=σ1σ2σ3+σ1ϵ0=(338+6)1038.851012=0.226109N/CE =\frac{\sigma_1- \sigma_2- \sigma_3+ \sigma_1 }{\epsilon_0}=\frac{(3-3-8+6)*10^{3}}{8.85*10^{-12}} = 0.226*10^{9} N/C